# Scientific Large Language Models: A Survey on Biological & Chemical Domains

### v1

@@ 7.2. Datasets
Dataset Time Scale Keywords
@@
PubChemSTM (Liu et al., 2023a) 2022.12 281K+250K Structure-Text Pairing
@@
Prot.&Text SwissProtCLAP (Liu et al., 2023f) 2023.02 441K Protein Sequence Analysis
ProtDescribe (Xu et al., 2023b) 2023.07 553K Protein Sequence Representation
@@
InstructProtein (Wang et al., 2023c) 2023.10 2.8M Structured KG, Bidirectional Protein-Text Generation
@@ 7.2.1. Molecule-Text Datasets
The molecule-text datasets are mainly sourced from chemical literature, patents, drug instructions, etc., which contain descriptions of molecules. After pre-processing (e.g., data clearing), it can be used to train the model to achieve communication between molecular language and natural language. .
@@
  • PubChemSTM (Liu et al., 2023a). PubChemSTM dataset is constructed using data from the PubChem (Kim et al., 2021a) database, which consists of 250,000 molecules and 281,000 structure-text pairs. PubChemSTM is used to train the MoleculeSTM model (Liu et al., 2023a), and can be extended to structure-text retrieval tasks and text-based molecule editing tasks.
@@ 7.2.2. Protein-Text Datasets
  • SwissProtCLAP (Liu et al., 2023f). The SwissProtCLAP dataset, derived from the SwissProt database (Boutet et al., 2007)—a subset of UniProt (Consortium, 2021), consists of 441,000 text-protein sequence pairs. This dataset finds extensive applications in bioinformatics and molecular biology, particularly in tasks requiring the integration of textual descriptions with protein sequences.
  • ProtDescribe (Xu et al., 2023b). The ProtDescribe dataset, comprising 553,052 aligned pairs of protein sequences and corresponding property descriptions, is constructed using data from the Swiss-Prot database (Bairoch and Apweiler, 2000). This database is renowned for its high-quality annotations of diverse protein properties. The dataset facilitates the incorporation of protein property information into protein language models, thereby enhancing their ability to model semantic correlations between known and unknown protein properties.
@@
  • InstructProtein (Wang et al., 2023c). The InstructProtein dataset is composed of protein-text instruction data derived from the protein knowledge graph constructed by the UniProt/Swiss-Prot database (Consortium, 2019), which encompasses a comprehensive collection of 2.8 million data. It is the first high-quality protein instruction dataset for tuning LLMs.
@@ 7.2.4. Comprehensive Datasets
  • Scientific Knowledge Dataset (Xie et al., 2023). The Scientific Knowledge Dataset is constructed using two main sources: the SciQ dataset (Welbl et al., 2017), which includes 13,679 science exam questions and answers, and a dataset of 6,057,261 full-text scholarly papers from the Web of Science. This comprehensive dataset is designed to train and enhance large language models for scientific tasks, such as question-answering and knowledge modeling in natural science domains.
@@ 7.3. Evaluation
Cross-modal Prediction. Cross-modal prediction is a task that involves using multimodal models to predict the functionality of molecules, proteins, and genomes based on textual instructions. These approaches combine different types of data, such as text and molecular structures, to make accurate predictions about the function of biological entities. The specific functions can be found in Section 4.3, 5.3, 6.3. Most encoder-based multimodal models such as MoleculeSTM (Liu et al., 2023a) are designed for this task. Moreover, decode-only architectures such as Mol-Instructions (Fang et al., 2023a) also possess the capability to directly generate functions. As a fundamental task, cross-modal function prediction has various applications in bioinformatics and drug discovery, including identifying potential drug targets, understanding the interactions between proteins, and predicting the activity of molecules.
@@
Cross-modal Retrieval. Cross-modal retrieval refers to the process of retrieving information from one modality based on a query from another modality, and vice versa. In the context of multimodal scientific language models, cross-modal retrieval primarily involves retrieving molecules, proteins, and genomes that are related to textual descriptions, and vice versa. For example, given a textual description of a molecule, a multimodal model can be used to retrieve similar molecules based on their functional descriptions. This can be useful in drug discovery, where researchers may want to find molecules that have specific properties. Models can accomplish these tasks include KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022) and MolFM (Luo et al., 2023b). On the other hand, given a protein or genomic dataset, a multimodal model such as ProtST-ESM-1b (Xu et al., 2023b) can be used to retrieve textual descriptions that provide information about the properties, functions, or interactions of the proteins or genes in the dataset. This can aid in understanding the biological mechanisms or relationships between different proteins or genes. Additionally, multimodal models can retrieve relevant protein sequences corresponding to a given molecular structure, or find the appropriate molecular structures based on a protein sequence. For example, DrugCLIP (Gao et al., 2023b) uses a dense retrieval approach, employing contrastive learning to maximize similarity between a protein-molecule pair, which enables two-way screening.
@@
Cross-modal Generation. Cross-modal generation tasks aim to generate data in one modality based on data from another modality, which involves establishing interconnections between the different modalities. MM-Sci-LLMs are designed to process data from multiple modalities including texts, molecules, proteins, and genomes. For example, Text2Molecule refers to generating molecular information (e.g., SMILES), based on input text descriptions. Many models can accomplish this task including Text2Mol (Edwards et al., 2021), MolT5 (Edwards et al., 2022), and MOL-Instruction (Fang et al., 2023a). The bidirectional conversion capability of cross-modal generation models allows them to not only generate data in one modality based on data from another modality but also to convert data from one modality to another. For example, Molecule2Text is to produce precise and pertinent textual descriptions derived from molecular data, namely molecule captioning. Similar to the mutual generation of molecule and text, Text2Protein aims to generate protein-related information including sequences and structures based on text descriptions. A representative model is ProteinDT (Liu et al., 2023f) which introduced a text-guided protein design method. Conversely, Protein2Text converts protein sequences into detailed text descriptions to express the complex biological characteristics of proteins. Prot2Text (Abdine et al., 2023), BioMedGPT-10B (Luo et al., 2023c) and MOL-Instruction (Fang et al., 2023a) can address this task effectively. Additionally, Protein2Molecule involves the process of generating relevant molecules from protein sequences, which can be achieved by ChemBERTaLM (Uludoğan et al., 2022) and DrugGPT (Li et al., 2023a). Overall, cross-modal generation in multi-modal models plays a crucial role in bridging the gap between different modalities and enabling the creation of new and cohesive multi-modal data.
@@ 8.1.2. Architectures and Learning Objectives
  • Autoregressive Learning Objective Limitations. Natural languages are generally processed in an autoregressive manner. However, this is not the case with biological and chemical languages, which are not inherently created or interpreted autoregressively. In these domains, sequences function as a whole. Therefore, an ideal learning objective for a Sci-LLM should allow the model to capture semantic information from the entire sequence, considering bidirectional context, rather than relying solely on one-way information flow.
@@ 8.1.3. Model Evaluation
However, this methodology predominantly addresses textual outputs. The challenge intensifies when evaluating the generation of biological and chemical languages, such as those used in the de novo design of proteins or molecules. Currently, it is arduous for computational systems, and even more so for human experts like biologists and chemists, to accurately ascertain the characteristics and efficacy of these generated entities based solely on observation. Computational methods like molecular simulation offer some insights, but their reliability can vary, sometimes deviating significantly from actual outcomes, not to mention that they are usually too slow to closely collaborate with Sci-LLMs.
The computational metrics discussed in Sec.4, Sec.5, and Sec. 6 offer indirect indications of the quality of newly generated proteins or molecules. However, these indicators are not always definitive. Wet-lab experiments remain the gold standard for validation, yet they are often beyond the scope and capabilities of many AI research teams. This gap underscores the urgent need for more robust and reliable computational evaluation and benchmarking systems. Such systems would not only augment the accuracy of assessments in the absence of wet-lab experimentation but also enhance the overall development and application of Sci-LLMs in scientific research.
@@ 8.2. Future Directions
  • Augmenting Sci-LLMs with Specialized Tools and Agents. General LLMs have demonstrated remarkable proficiency in leveraging software and tools, such as calculators, to accomplish complex tasks. This capability is equally pertinent to Sci-LLMs, which can be significantly enhanced through integration with science-specific computational tools. A prime example of this is the recent success of GPT-4, which has shown proficiency in web searching, academic paper analysis, and designing experiments, including replicating intricate cross-coupling reactions like the Suzuki and Sonogashira experiments. The realm of scientific tools is vast and intricate, encompassing a wide array of computational resources. The challenge of accurately selecting the appropriate tool or software from tens of thousands of possibilities, precisely inputting data, and effectively interpreting the output, presents a rich area for research exploration. The advancement in this area holds significant promise for augmenting the capabilities of Sci-LLMs, enabling them to perform more sophisticated and specialized scientific tasks with higher precision and efficiency.
@@
  • Sec.3: Jing Yu, Keyan Ding, Qiang Zhang, Zhuoyi Xiang, Huajun Chen
@@
  • Sec.5: Yiwen Zhang, Keyan Ding, Qiang Zhang, Zeyuan Wang, Hongyang Chen, Huajun Chen

### v2

@@ 7.2. Datasets
Dataset Last updated Scale Keywords
@@
PubChemSTM (Liu et al., 2023b) 2022.12 281K+250K Structure-Text Pairing
@@
MoleculeQA (Lu et al., 2024) 2024.03 62K+23K Comprehensive Evaluation of Molecular
Prot.&Text SwissProtCLAP (Liu et al., 2023g) 2023.02 441K Protein Sequence Analysis
ProtDescribe (Xu et al., 2023c) 2023.07 553K Protein Sequence Representation
@@
InstructProtein (Wang et al., 2023g) 2023.10 2.8M Structured KG, Bidirectional Protein-Text Generation
ProteinLMDataset (Shen et al., 2024) 2024.06 89.3K Protein Sequence-Text
ProteinLMBench (Shen et al., 2024) 2024.06 944 A benchmark for protein understanding
Cell&Text GEO (Clough and Barrett, 2016) 2016 1.3M High-throughput Genomics
HCA (Regev et al., 2017) 2017.12 - Single-cell Transcriptomics
ARCHS4 (Lachmann et al., 2018) 2018.04 188K Mouse and human RNA-seq samples
NCBI (Wheeler et al., 2007) 2020.10 3B Comprehensive Biotechnological Data
cellxgene (Megill et al., 2021) 2021.04 37M High-dimensional sparse matrices
SRT (He et al., 2021a) 2021.11 2.7M Spatial transcriptomic data
CellTypist (Domínguez Conde et al., 2022) 2022.05 360K Immune cell analysis, Cell type annotation
scLibrary (Zhao et al., 2024b) 2024.06 27.5M Extract from cellxgene
@@
SMolInstruct (Yu et al., 2024) 2024.02 3.3M Comprehensive Instruction, Diverse Text Descriptions
@@ 7.2.1. Molecule-Text Datasets
The molecule-text datasets are mainly sourced from chemical literature, patents, drug instructions, etc., which contain descriptions of molecules. After pre-processing (e.g., data clearing), it can be used to train the model to achieve communication between molecular language and natural language.
@@
  • PubChemSTM (Liu et al., 2023b). PubChemSTM dataset is constructed using data from the PubChem (Kim et al., 2021a) database, which consists of 250,000 molecules and 281,000 structure-text pairs. PubChemSTM is used to train the MoleculeSTM model (Liu et al., 2023b), and can be extended to structure-text retrieval tasks and text-based molecule editing tasks.
@@
  • MoleculeQA (Lu et al., 2024). MoleculeQA is a comprehensive question-answering dataset consisting of 62K pairs that evaluate factual accuracy for molecular comprehension. It covers a range of molecules with detailed descriptions linked to authoritative sources, providing a robust platform for assessing language models’ capabilities in understanding molecular properties, sources, structures, and applications.
@@ 7.2.2. Protein-Text Datasets
  • SwissProtCLAP (Liu et al., 2023g). The SwissProtCLAP dataset, derived from the SwissProt database (Boutet et al., 2007)—a subset of UniProt (Consortium, 2021), consists of 441,000 text-protein sequence pairs. This dataset finds extensive applications in bioinformatics and molecular biology, particularly in tasks requiring the integration of textual descriptions with protein sequences.
  • ProtDescribe (Xu et al., 2023c). The ProtDescribe dataset, comprising 553,052 aligned pairs of protein sequences and corresponding property descriptions, is constructed using data from the Swiss-Prot database (Bairoch and Apweiler, 2000). This database is renowned for its high-quality annotations of diverse protein properties. The dataset facilitates the incorporation of protein property information into protein language models, thereby enhancing their ability to model semantic correlations between known and unknown protein properties.
@@
  • InstructProtein (Wang et al., 2023g). The InstructProtein dataset is composed of protein-text instruction data derived from the protein knowledge graph constructed by the UniProt/Swiss-Prot database (Consortium, 2019), which encompasses a comprehensive collection of 2.8 million data. It is the first high-quality protein instruction dataset for tuning LLMs.
  • ProteinLMDataset (Shen et al., 2024). The InstructProtein dataset is composed of protein-text instruction data derived from the protein knowledge graph constructed by the UniProt/Swiss-Prot database (Consortium, 2019), which encompasses a comprehensive collection of 2.8 million data. It is the first high-quality protein instruction dataset for tuning LLMs.
  • ProteinLMBench (Shen et al., 2024). ProteinLMBench stands as the inaugural benchmark for evaluating the protein comprehension proficiency of LLMs. It encompasses 944 meticulously curated multiple-choice questions that span a spectrum of protein-related tasks. This benchmark is designed to rigorously assess the interpretative and analytical skills of LLMs in the context of protein sequence understanding
7.2.3. Gene/Cell-Text
The integration of single-cell transcriptomics and natural language text data transforms gene expression profiles into textual sequences, enabling the application of large language models to biological data. This approach facilitates biologically relevant cell generation, accurate cell type predictions, and meaningful biological insights extraction through natural language processing.
  • GEO (Clough and Barrett, 2016). The Gene Expression Omnibus (GEO) is a vast public repository that archives high-throughput gene expression and functional genomics data sets, accommodating various technologies such as DNA microarrays, high-throughput sequencing, and genome methylation studies. With over 54,640 studies and more than 1.3 million samples from 2,889 different organisms, GEO provides extensive tools for data retrieval, visualization, and analysis, supporting researchers globally in exploring and interpreting complex genomic data.
  • HCA (Regev et al., 2017). The Human Cell Atlas dataset is a comprehensive reference map of all human cell types, characterized by high-throughput single-cell transcriptomics. It aims to profile millions of individual cells to understand their molecular signatures, locations, and interactions within healthy human tissues. This dataset provides invaluable insights into human physiology, development, and disease mechanisms.
  • ARCHS4 (Lachmann et al., 2018). The dataset described is a comprehensive RNA-seq collection encompassing 187,946 samples, with 103,083 from mice and 84,863 from humans. This large-scale dataset is processed and accessible through the ARCHS4 web resource, offering extensive tools for gene expression analysis and visualization. It provides a valuable resource for retrospective and integrative analyses in genomic research.
  • NCBI (Wheeler et al., 2007). The National Center for Biotechnology Information (NCBI) offers a comprehensive suite of biological databases, including GenBank for nucleic acid sequences and PubMed for life science literature. The Entrez system provides integrated access to 34 distinct databases, collectively containing billions of records across genomics, genetics, proteins, chemicals, and more. This extensive resource supports a wide range of biological research and data retrieval needs.
  • cellxgene (Megill et al., 2021). The cellxgene dataset in question consists of high-dimensional sparse matrices, commonly used in single-cell RNA sequencing (scRNAseq) analyses. These matrices typically include millions of individual cell observations and tens of thousands of gene expression measurements, enabling detailed exploration and analysis of cellular heterogeneity and gene expression patterns across different conditions and populations.
  • SRT (He et al., 2021a). The Spatial Transcriptomics (SRT) data set, specifically designed for high-plex molecular imaging, provides comprehensive insights into gene expression and protein localization within tissue sections. This data set allows for the spatial mapping of RNA and protein targets on FFPE (formalin-fixed, paraffin-embedded) samples, enabling detailed analysis at single-cell and subcellular resolutions. The utilization of advanced imaging technologies facilitates a deeper understanding of tissue architecture and cellular interactions in various biological contexts.
  • CellTypist (Domínguez Conde et al., 2022). The CellTypist dataset comprises nearly 360,000 single-cell transcriptomes obtained from 16 different tissues of 12 deceased individuals. It was generated using single-cell RNA sequencing (scRNA-seq) and VDJ sequencing, focusing on immune cell profiling. The dataset supports comprehensive cell type annotation and the development of the CellTypist tool, providing detailed insights into the heterogeneity of immune cells across human tissues.
  • scLibrary (Zhao et al., 2024b). The scLibrary dataset, constructed from cellxgene, includes approximately 27.5 million pairs of single-cell RNA sequencing (scRNA-seq) data and their corresponding textual descriptions. It features comprehensive metadata annotations from various biological ontologies, providing detailed information on cell type, developmental stage, tissue, disease status, and more. This dataset enables enhanced understanding of cell identity through integrated transcriptomic and textual data.
@@ 7.2.5. Comprehensive Datasets
  • Scientific Knowledge Dataset (Xie et al., 2023). It is constructed using two main sources: the SciQ dataset (Welbl et al., 2017), which includes 13,679 science exam questions and answers, and a dataset of 6,057,261 full-text scholarly papers from the Web of Science. This comprehensive dataset is designed to train and enhance large language models for scientific tasks, such as question-answering and knowledge modeling in natural science domains.
@@
  • SMolInstruct (Yu et al., 2024). It is a large-scale and high-quality instruction tuning dataset designed for advancing large language models in the field of chemistry. It encompasses 14 distinct chemistry tasks with over three million samples, facilitating capabilities such as name conversion, property prediction, molecule description, and chemical reaction analysis. This dataset serves as a robust foundation for training and evaluating LLMs on a wide array of chemistry-related tasks, pushing the boundaries of their performance in this scientific domain.
@@ 7.3. Evaluation
Cross-modal Prediction. Cross-modal prediction is a task that involves using multimodal models to predict the functionality of molecules, proteins, and genomes based on textual instructions. These approaches combine different types of data, such as text and molecular structures, to make accurate predictions about the function of biological entities. The specific functions can be found in Section 4.3, 5.3, 6.3. Most encoder-based multimodal models such as MoleculeSTM (Liu et al., 2023b) are designed for this task. Moreover, decode-only architectures such as Mol-Instructions (Fang et al., 2023a) also possess the capability to directly generate functions. As a fundamental task, cross-modal function prediction has various applications in bioinformatics and drug discovery, including identifying potential drug targets, understanding the interactions between proteins, and predicting the activity of molecules.
@@
Cross-modal Retrieval. Cross-modal retrieval refers to the process of retrieving information from one modality based on a query from another modality, and vice versa. In the context of multimodal scientific language models, cross-modal retrieval primarily involves retrieving molecules, proteins, and genomes that are related to textual descriptions, and vice versa. For example, given a textual description of a molecule, a multimodal model can be used to retrieve similar molecules based on their functional descriptions. This can be useful in drug discovery, where researchers may want to find molecules that have specific properties. Models can accomplish these tasks include KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022) and MolFM (Luo et al., 2023b). On the other hand, given a protein or genomic dataset, a multimodal model such as ProtST-ESM-1b (Xu et al., 2023c) can be used to retrieve textual descriptions that provide information about the properties, functions, or interactions of the proteins or genes in the dataset. This can aid in understanding the biological mechanisms or relationships between different proteins or genes. Additionally, multimodal models can retrieve relevant protein sequences corresponding to a given molecular structure, or find the appropriate molecular structures based on a protein sequence. For example, DrugCLIP (Gao et al., 2023b) uses a dense retrieval approach, employing contrastive learning to maximize similarity between a protein-molecule pair, which enables two-way screening.
@@
Cross-modal Generation. This tasks aim to generate data in one modality based on another modality. MM-Sci-LLMs are designed to process data from multiple modalities including texts, molecules, proteins, and genomes. For example, Text2Molecule refers to generating molecular information (e.g., SMILES), based on input text descriptions. Many models can accomplish this task including Text2Mol (Edwards et al., 2021), MolT5 (Edwards et al., 2022), and MOL-Instruction (Fang et al., 2023a). The bidirectional conversion capability of cross-modal generation models allows them to not only generate data in one modality based on data from another modality but also to convert data from one modality to another. For example, Molecule2Text is to produce precise and pertinent textual descriptions derived from molecular data, namely molecule captioning. Similar to the mutual generation of molecule and text, Text2Protein aims to generate protein-related information including sequences and structures based on text descriptions. A representative model is ProteinDT (Liu et al., 2023g) which introduced a text-guided protein design method. Conversely, Protein2Text converts protein sequences into detailed text descriptions to express the complex biological characteristics of proteins. Prot2Text (Abdine et al., 2023), BioMedGPT-10B (Luo et al., 2023c) and MOL-Instruction (Fang et al., 2023a) can address this task effectively. Additionally, Protein2Molecule involves the process of generating relevant molecules from protein sequences, which can be achieved by ChemBERTaLM (Uludoğan et al., 2022) and DrugGPT (Li et al., 2023a). Overall, cross-modal generation in multi-modal models plays a crucial role in bridging the gap between different modalities and enabling the creation of new and cohesive multi-modal data.
@@ 8.1.2. Architectures and Learning Objectives
  • Autoregressive Learning Objective Limitations. Natural language generation is generally processed in an autoregressive manner. However, this is not the case with biological and chemical languages, which are not inherently created or interpreted autoregressively. In these domains, sequences function as a whole. Therefore, an ideal learning objective for a generative Sci-LLM should allow the model to capture semantic information from the entire sequence, considering bidirectional context, rather than relying solely on one-way information flow.
@@ 8.1.3. Model Evaluation
However, this methodology predominantly addresses textual outputs. The challenge intensifies when evaluating the generation of biological and chemical languages, such as those used in the de novo design of proteins or molecules. Currently, it is arduous for computational systems, and even more so for human experts like biologists and chemists, to accurately ascertain the characteristics and efficacy of these generated entities based solely on observation. Computational methods like molecular simulation offer some insights, but their reliability can vary, sometimes deviating significantly from actual outcomes, not to mention that they are usually too slow to closely collaborate with Sci-LLMs. The computational metrics discussed in Sec.4, Sec.5, and Sec. 6 offer indirect indications of the quality of newly generated proteins or molecules. However, these indicators are not always definitive. Wet-lab experiments remain the gold standard for validation, yet they are often beyond the scope and capabilities of many AI research teams. This gap underscores the urgent need for more robust and reliable computational evaluation and benchmarking systems. Such systems would not only augment the accuracy of assessments in the absence of wet-lab experimentation but also enhance the overall development and application of Sci-LLMs in scientific research.
@@ 8.2. Future Directions
  • Augmenting Sci-LLMs with Specialized Tools and Agents. General LLMs have demonstrated remarkable proficiency in leveraging software and tools, such as calculators, to accomplish complex tasks. This capability is equally pertinent to Sci-LLMs, which can be significantly enhanced through integration with science-specific computational tools (Ramos et al., 2024). A prime example of this is the recent success of GPT-4, which has shown proficiency in web searching, academic paper analysis, and designing experiments, including replicating intricate cross-coupling reactions like the Suzuki and Sonogashira experiments. The realm of scientific tools is vast and intricate, encompassing a wide array of computational resources. The challenge of accurately selecting the appropriate tool or software from tens of thousands of possibilities, precisely inputting data, and effectively interpreting the output, presents a rich area for research exploration. The advancement in this area holds significant promise for augmenting the capabilities of Sci-LLMs, enabling them to perform more sophisticated and specialized scientific tasks with higher precision and efficiency.
@@
  • Sec.3: Jing Yu, Keyan Ding, Qiang Zhang, Kehua Feng, Tao Huang, Pengju Yan, Xiaolin Li, Huajun Chen
@@
  • Sec.5: Yiwen Zhang, Keyan Ding, Qiang Zhang, Zeyuan Wang, Zhuoyi Xiang, Hongyang Chen, Huajun Chen